OpenStudy (anonymous):

1. 4632.73 = 2000e^0.12(t) how do i solve for t? 2. 100e^0.05(t)=10e^0.15(t) solve for t

OpenStudy (anonymous):

e^.12t=4632.73/2000=2.316365 take log .12t= ln(2.316365)=0.839999147 t=0.839999147/.12=7

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

m not sir...

OpenStudy (anonymous):

opps sorry so for the other problems should i divide 10 first and take the natural log on both side?

OpenStudy (anonymous):

100e^0.05(t)=10e^0.15(t) 100/10=e^(.10(t)) now proceed ..

OpenStudy (anonymous):

t = 23?

OpenStudy (anonymous):

yes it may...

OpenStudy (anonymous):

so you just subtract e^0.05 from e^0.15 is that the rules?